# Robotic Manipulation via Imitation Learning: Taxonomy, Evolution, Benchmark, and Challenges

### v1

@@ Abstract
Robotic manipulation (RM) is central to enabling autonomous robots to interact with and alter their environments in real-world scenarios. Among the learning paradigms, imitation learning has emerged as a powerful approach, allowing robots to rapidly acquire complex manipulation skills from human demonstrations. This survey provides the first systematic review dedicated to imitation learning for robotic manipulation. We identify and analyze a large set of representative studies selected for their scientific quality and community impact. For each, we provide a structured summary covering purpose, technical implementation, taxonomy, input formats, priors, strengths, limitations, and citation metrics. Beyond cataloging, we trace the chronological evolution of imitation learning techniques within robotic manipulation policies (RMPs), highlighting key methodological shifts—from diffusion and flow matching to autoregressive and affordance-driven strategies. Where available, we compile benchmark results and conduct quantitative comparisons, enabling an integrated view of performance across tasks and environments. Finally, we outline open challenges such as generalization, embodiment diversity, data efficiency, and benchmark standardization, and we discuss promising directions toward scalable and general-purpose robotic manipulation. By synthesizing methods, benchmarks, and challenges, this survey aims to serve both as an entry point for newcomers and a reference for active researchers seeking to advance imitation learning in robotic manipulation.
@@ I-A Motivation
Robotic manipulation (RM) refers to the ability of robots to physically interact with and transform their surroundings by grasping, moving, assembling, or otherwise altering objects. It is a fundamental capability of autonomous systems and underpins their deployment in real world. The importance of manipulation has long been recognized: already Aristotle described the human hand as “the tool of tools” [1], while Anaxagoras argued that “man is the most intelligent of the animals because he has hands” [2]. RM is crucial because it enables automation of tasks that are too dangerous, precise, repetitive, or labor-intensive for humans, spanning domains such as manufacturing, healthcare, logistics, and household assistance. By extending human reach into hazardous or delicate settings, RM enhances safety, efficiency, and productivity. Achieving robust RM, however, requires the development of accurate and adaptive control policies that allow robots to act autonomously in dynamic and unstructured environments—an enduring challenge at the core of modern robotics.
A robotic manipulation policy (RMP) specifies how a robot selects and executes actions based on its sensory observations to achieve a manipulation goal. However, the space of possible object states, contact dynamics, and task variations is combinatorially large, making it impossible to encode effective strategies through explicit rules or heuristics. Consequently, current state-of-the-art approaches are predominantly data-driven, leveraging deep learning to automatically learn robust representations and control strategies from demonstrations or large-scale datasets. Yet, even these approaches face significant challenges, since robots must operate in dynamic and unstructured environments, where not only object properties and task goals but also environmental conditions—such as lighting, clutter, occlusions, or background changes—can vary unpredictably. To address these difficulties, imitation learning has emerged as a powerful paradigm, enabling robots to acquire complex skills directly from human demonstrations. Recent progress—amplified by advances in computer vision and large language models (LLMs)—has further enhanced robots’ ability to perceive, reason, and plan actions. This survey focuses on the role of imitation learning for RMPs, providing a comprehensive analysis of key methodologies, benchmarks, and applications, and outlining future directions toward scalable and general-purpose RM.
@@
In light of the fast-paced developments in RMP, this review primarily considers research conducted between 2021 and 2025. Our motivation is to provide newcomers to this field with a rapid and comprehensive understanding of the current research landscape, while also supporting active researchers by helping them efficiently locate related works of interest. By presenting the hierarchical taxonomy, purpose, input, and pre-training data formats, strengths, limitations, and citation of each paper, we aim to distill the core information and enable researchers to quickly assess which approaches align with their specific needs or application scenarios. Furthermore, we present quantitative comparisons across different benchmarks. Additionally, we summarize highly sought-after application areas and key existing challenges in the field, intending to mobilize collective efforts within the community to accelerate progress.
@@ II-A1 Continuous Action
Flow-matching, as a novel generative model, has garnered significant attention due to its ability to generate high-quality results more quickly compared to diffusion models. Many researchers have explored its potential in combination with robotic action generation, leading to the emergence of several outstanding works in this area. Initially, FMP [47] gave the first attempt to ground vision-language model (VLM) affordance with flow matching for robot manipulation, enabling more stable training and faster inference. Reactive diffusion policy (RDP) [48], a novel slow-fast visual-tactile imitation learning algorithm, allowing robots to adjust their strategies in real-time for contact-rich manipulation tasks. ActionFlow [49] introduces an $\mathrm{SE}$(3) invariant Transformer, which enables informed spatial reasoning based on the relative $\mathrm{SE}$(3) poses between observations and actions. Later, $\pi_{0}$ [50] integrates pre-trained VLM with flow matching to inherit Internet-scale semantic knowledge, enhancing the robot’s ability to efficiently learn and execute a wide variety of tasks, such as laundry folding, table cleaning, and assembling boxes. GraspVLA [51] integrates autoregressive perception tasks and flow- matching-based action generation into a unified chain-of-thought process, enabling joint training on synthetic action data and Internet semantics data, exhibiting direct sim-to-real transfer and strong zero-shot generalization, as well as few-shot adaptability to specialized scenarios and human preferences. This was further advanced by HiRobot [52], which utilizes VLMs for both high-level reasoning and low-level task execution, enabling the processing of much more complex prompts. Most recently, SmolVLA [53] focused on making these techniques more efficient in resource-constrained environments, incorporating lightweight visual learning systems while maintaining the robustness of flow matching. Overall, these developments illustrate the continuous refinement of flow-matching techniques, expanding their applicability from basic manipulation to more complex, adaptable, and resource-efficient systems.
@@ II-A2 Discrete Action
Meanwhile, because robot demonstration data are scarce, several works have begun exploring how to learn useful action representations from large-scale actionless videos. LAPA [80] proposes unsupervised latent-action pretraining from actionless videos by learning discrete latent actions and then finetuning to map latents to robot actions, enabling large-scale pretraining without costly robot action labels. UniVLA [79] learns task-centric latent action representations from heterogeneous, cross-embodiment videos so a single policy can be decoded onto different robots, enabling scalable, compute-efficient pretraining that transfers across embodiments and environments. WorldVLA [82] is the first to unifie an discrete autoregressive world model and an action model in one VLA framework, so images and actions are jointly modeled. The coupled model improves video prediction and action performance, and introduces masking attention strategies to reduce autoregressive action error accumulation.
@@ III-C Policy for Multi Scenario
A few years back, Gato [60] initiated the idea of creating a multi-purpose “generalist agent”, capable of solving a set of tasks, even tasks it has never seen. A main issue to solve this problem is access to reliable data. Recently, the scale of robotic data has increased by a large margin thanks to community initiatives such as Open-X [106]. This permitted new models to emerge by getting inspiration from other fields such as natural language processing and computer vision for pretraining (see Section IV). Leveraging the reasoning capacities of pretrained Visual Language Models, a set of new methods was introduced to obtain agents capable of understanding and reasoning over scenes before acting, leading to multi-scenario policies capable of generalizing. One of the first iteration of such model is RT-1 [103] which pretrains a visual model with text conditioning on a large-scale dataset of tasks. Following work leverages pretrained models on out-of-domain data for their strong reasoning capabilities to learn an action model, which leads to the set of VLAs;
@@ III-F Policy for Generalization
For embodiment generalization, Cogact [40], UniVLA [79], and SpatialVLA [74] provide examples of good generalization to unseen embodiments and tasks. Hamster [76], TraceVLA [71], and OpenVLA [67] exhibit strong generalization to different embodiments and scenes, accommodating varying lighting and distractors. Despite these initial approaches, generalization across different embodiments is still limited to similar morphologies and application scenarios. Currently, the generalization of manipulation strategies between embodiments with different configurations, particularly those with varying degrees of freedom, has not yet been achieved.
@@ III-H Policy for Sampling Efficiency
Other works improve efficiency by replacing single-action generation with action trunk, such as ACT [63], VQ-BeT [66], and ARP [69]. Transformer-based methods also benefit from more efficient action token usage: Fast [73] introduces a frequency-space action sequence tokenization that greatly compresses token count, while VLACache [75] selects and reuses tokens with minimal changes across steps. CARP [70] adopts a coarse-to-fine autoregressive strategy, first learning multi-scale action representations and then refining them via a GPT-style transformer. DiVLA [39] and RDP [48] enhance efficiency by generating shared action tokens in latent space.
@@ VII-B Summary of open challenges
The wide variety of robot designs and configurations leads to significant differences in the input and output spaces for learning manipulation policies. These discrepancies make it difficult for existing models to generalize across different robotic platforms. As a result, a unified manipulation strategy that can work across various robot types is challenging to achieve. To address this, having a standardized robot configuration is essential for building a versatile and practical foundational VLA model. This would allow for more effective and transferable learning across different robotic systems and tasks.

### v2

@@ Abstract
Robotic manipulation (RM) is central to enabling autonomous robots to interact with and alter their environments in real-world scenarios. Among the learning paradigms, imitation learning has emerged as a powerful approach, allowing robots to rapidly acquire complex manipulation skills from human demonstrations. This survey provides the first systematic review dedicated to imitation learning for robotic manipulation. We identify and analyze a large set of representative studies selected for their scientific quality and community impact. For each, we provide a structured summary covering purpose, technical implementation, taxonomy, input formats, priors, strengths, limitations, and citation metrics. Beyond cataloging, we trace the chronological evolution of imitation learning techniques within robotic manipulation policies (RMPs), highlighting key methodological shifts—from diffusion and flow matching to autoregressive and affordance-driven policies. Where available, we compile benchmark results and conduct quantitative comparisons, enabling an integrated view of performance across tasks and environments. Finally, we outline open challenges such as generalization, embodiment diversity, data efficiency, and benchmark standardization, and we discuss promising directions toward scalable and general-purpose RMPs. By synthesizing methods, benchmarks, and challenges, this survey aims to serve both as an entry point for newcomers and a reference for active researchers seeking to advance imitation learning in RMPs.
@@ I-A Motivation
Robotic manipulation (RM) refers to the ability of robots to physically interact with and transform their surroundings by grasping, moving, assembling, or otherwise altering objects. It is a core capability for deploying autonomous systems in the real world. The importance of manipulation has long been recognized: already Aristotle described the human hand as “the tool of tools” [1], while Anaxagoras argued that “man is the most intelligent of the animals because he has hands” [2]. RM is crucial because it enables automation of tasks that are too dangerous, precise, repetitive, or labor-intensive for humans, spanning domains such as manufacturing, healthcare, logistics, and household assistance. By extending human reach into hazardous or delicate settings, RM enhances safety, efficiency, and productivity. Achieving robust RM, however, requires the development of robust control policies that allow robots to act adaptively in dynamic and unstructured environments—an enduring challenge at the core of modern robotics.
A robotic manipulation policy (RMP) specifies how a robot selects and executes actions based on its sensory observations to achieve a manipulation goal, i.e., $\pi:\mathcal{O}\times\mathcal{G}\to\mathcal{A},a_{t}=\pi(o_{t},g)$. Here $\mathcal{O}$ is the observation space, $\mathcal{G}$ is the goal space (e.g., language instructions, object states), and $\mathcal{A}$ is the action space. However, the space of possible object states, contact dynamics, and task variations is combinatorially large, making it impossible to encode effective policies through explicit rules or heuristics. Consequently, current state-of-the-art approaches are predominantly data-driven, leveraging deep learning to learn representations and control policies from large-scale datasets. Yet, even these approaches face significant challenges, since robots must operate in dynamic and unstructured environments, where not only object properties and task goals but also environmental conditions—such as lighting, clutter, occlusions, or background changes—can vary unpredictably. To address these difficulties, imitation learning (IL) has emerged as a powerful paradigm that learns a policy from expert demonstrations $\mathcal{D}=\{\{(o_{t}^{i},a_{t}^{i})\}_{t=1}^{T_{i}}\}_{i=1}^{N}$, with the objective $\pi_{\theta}\approx\pi_{E}$, where $\pi_{E}$ is the expert policy. Recent progress amplified by advances in computer vision and large language models (LLMs) has further enhanced robots’ ability to perceive, reason, and plan actions. This survey focuses on IL-based RMPs, providing a comprehensive analysis of methodologies, benchmarks, applications, and outlining future directions toward scalable and general-purpose RMPs.
@@
In light of the fast-paced developments in RMP, this review primarily considers research conducted between 2021 and 2025. Our motivation is to provide newcomers to this field with a rapid and comprehensive understanding of the current research landscape, while also supporting active researchers by helping them efficiently locate related works of interest. By presenting the hierarchical taxonomy, purpose, input, and pre-training strategies, strengths, limitations, and citations of each paper, we aim to distill the core information and enable researchers to quickly assess which approaches align with their specific needs. Furthermore, we present quantitative comparisons across different benchmarks and summarize highly sought-after application areas and key existing challenges in the field, intending to mobilize collective efforts within the community to accelerate progress.
@@ II-A1 Continuous Action
Flow-matching, as a novel generative model, has garnered significant attention due to its ability to generate high-quality results more quickly compared to diffusion models. Many researchers have explored its potential in combination with robotic action generation, leading to the emergence of several outstanding works in this area. Initially, FMP [47] gave the first attempt to ground vision-language model (VLM) affordance with flow matching for robot manipulation, enabling more stable training and faster inference. Reactive diffusion policy (RDP) [48], a novel slow-fast visual-tactile imitation learning algorithm, allowing robots to adjust their policies in real-time for contact-rich manipulation tasks. ActionFlow [49] introduces an $\mathrm{SE}$(3) invariant Transformer, which enables informed spatial reasoning based on the relative $\mathrm{SE}$(3) poses between observations and actions. Later, $\pi_{0}$ [50] integrates pre-trained VLM with flow matching to inherit Internet-scale semantic knowledge, enhancing the robot’s ability to efficiently learn and execute a wide variety of tasks, such as laundry folding, table cleaning, and assembling boxes. GraspVLA [51] integrates autoregressive perception tasks and flow- matching-based action generation into a unified chain-of-thought process, enabling joint training on synthetic action data and Internet semantics data, exhibiting direct sim-to-real transfer and strong zero-shot generalization, as well as few-shot adaptability to specialized scenarios and human preferences. This was further advanced by HiRobot [52], which utilizes VLMs for both high-level reasoning and low-level task execution, enabling the processing of much more complex prompts. Most recently, SmolVLA [53] focused on making these techniques more efficient in resource-constrained environments, incorporating lightweight visual learning systems while maintaining the robustness of flow matching. Overall, these developments illustrate the continuous refinement of flow-matching techniques, expanding their applicability from basic manipulation to more complex, adaptable, and resource-efficient systems.
@@ II-A2 Discrete Action
Meanwhile, because robot demonstration data are scarce, several works have begun exploring how to learn useful action representations from large-scale actionless videos. LAPA [80] proposes unsupervised latent-action pretraining from actionless videos by learning discrete latent actions and then finetuning to map latents to robot actions, enabling large-scale pretraining without costly robot action labels. UniVLA [79] learns task-centric latent action representations from heterogeneous, cross-embodiment videos so a single policy can be decoded onto different robots, enabling scalable, compute-efficient pretraining that transfers across embodiments and environments. WorldVLA [82] unify a discrete autoregressive world model and an action model in one VLA framework, so images and actions are jointly modeled. The coupled model improves video prediction and action performance, and introduces masking attention strategies to reduce autoregressive action error accumulation.
@@ III-C Policy for Multi Scenario
A few years back, Gato [60] initiated the idea of creating a multi-purpose “generalist agent”, capable of solving a set of tasks, even tasks it has never seen. A main issue to solve this problem is access to reliable data. Recently, the scale of robotic data has increased by a large margin thanks to community initiatives such as Open-X [106]. This permitted new models to emerge by getting inspiration from other fields such as natural language processing and computer vision for pretraining (see Section IV). Leveraging the reasoning capacities of pretrained Visual Language Models, a set of new methods was introduced to obtain agents capable of understanding and reasoning over scenes before acting, leading to multi-scenario policies capable of generalizing. One of the first iteration of such model is RT-1 [103] which pretrains a visual model with text conditioning on a large-scale dataset of tasks. Following work leverages pretrained models on out-of-domain data for their strong reasoning capabilities to learn an action model, which leads to the set of VLAs.
@@ III-F Policy for Generalization
For embodiment generalization, Cogact [40], UniVLA [79], and SpatialVLA [74] provide examples of good generalization to unseen embodiments and tasks. Hamster [76], TraceVLA [71], and OpenVLA [67] exhibit strong generalization to different embodiments and scenes, accommodating varying lighting and distractors. Despite these initial approaches, generalization across different embodiments is still limited to similar morphologies and application scenarios. Currently, the generalization of manipulation policies between embodiments with different configurations, particularly those with varying degrees of freedom, has not yet been achieved.
@@ III-H Policy for Sampling Efficiency
Other works improve efficiency by replacing single-action generation with action trunk, such as ACT [63], VQ-BeT [66], and ARP [69]. Transformer-based methods also benefit from more efficient action token usage: Fast [73] introduces a frequency-space action sequence tokenization that greatly compresses token count, while VLACache [75] selects and reuses tokens with minimal changes across steps. CARP [70] adopts a coarse-to-fine autoregressive policy, first learning multi-scale action representations and then refining them via a GPT-style transformer. DiVLA [39] and RDP [48] enhance efficiency by generating shared action tokens in latent space.
@@ VII-B Summary of open challenges
The wide variety of robot designs and configurations leads to significant differences in the input and output spaces for learning manipulation policies. These discrepancies make it difficult for existing models to generalize across different robotic platforms. As a result, a unified manipulation policy that can work across various robot types is challenging to achieve. To address this, having a standardized robot configuration is essential for building a versatile and practical foundational VLA model. This would allow for more effective and transferable learning across different robotic systems and tasks.